OpenStudy (anonymous):

If 250 mg of a radioactive element decays to 240 mg in 12 hours, find the half-life of the element. is it setup like this: 240=250e^((-ln2/x)12)

OpenStudy (campbell_st):

well I would have said the equation was set up as \[240 = 250e^{12k}\] the general form is \[A = A_{0} e^{kt}\] k is the decay constant and t is time you will 1st need to find the decay constant... so \[\frac{240}{250} = e^{12k}\] take the base e log of both sides \[\ln(\frac{240}{250} )= 12k \] solve for k... and I'd say you will need 4 or 5 decimal places... and it should be negative.

OpenStudy (campbell_st):

for the time to have life \[125 = 250e^{kt}\] use the value of k from the 1st part and solve for t.

OpenStudy (anonymous):

why is it 125 and not 240

OpenStudy (anonymous):

i think he made a mistake.......

OpenStudy (campbell_st):

the initial quantity is 250 so half life is when it gets to 125.

OpenStudy (anonymous):

oh. it was the right answer. i just wasnt sure why you did that. so when its half life you do that. cool thanks